OpenStudy (anonymous):

exact value of cos (5pi/18-pi/9) ?

OpenStudy (anonymous):

\[\frac{5 \Pi}{18} - \frac{\Pi}{9} = ?\]

OpenStudy (anonymous):

5pi/18-2pi/18

OpenStudy (anonymous):

Good. Subtract and simplify.

OpenStudy (anonymous):

pi/6 ? I'm not sure how to find the exact value of this

OpenStudy (anonymous):

Very good! Cos(pi/6) is one of those values that you may or may not have been told to memorize at this point. If not, we can find it, as it's half the angle of pi/3. So, if you don't have it memorized, but do have cos(pi/3) memorized, we can use the half-angle formula to find it.

OpenStudy (anonymous):

Half-Angle Formula: \[Cos^2 (\frac{\Theta}{2}) = \frac{1}{2}(1+Cos(\Theta))\]

OpenStudy (anonymous):

cos 30 degrees? I believe it would be square root of 3 over 2

OpenStudy (anonymous):

That's the one :)

OpenStudy (anonymous):

If you already have that memorized, you're done. If not, you can show it by using the formula I posted, and Theta = pi/3

OpenStudy (anonymous):

Thank you so much that was very helpful

OpenStudy (anonymous):

Cos(pi/3) = 1/2 So: \[Cos^2(\frac{\Pi}{6}) = \frac{1}{2}(1+Cos(\frac{\Pi}{3}) = \frac{1}{2}(1 + \frac{1}{2})\] etc.

OpenStudy (anonymous):

My pleasure :)

OpenStudy (anonymous):

That makes sense now!